OpenStudy (anonymous):

Determine which equation below has a solution of x = 2. A. 5x + 8 – 12x = –6x + 2x – 3 B. 6x + 4 – 12x = –10x + 18x – 3 C. –4x – 6 + x = –2x + x – 10 D. –3x – 8 + x = –4x + x – 10

OpenStudy (jadeishere):

5x + 8 - 12x = -6x + 2x - 3 -7x + 8 = -4x - 3 (isolating the variable) -3x = -11 So not the first one Just keep cancaling out the ones that won't work

myininaya (myininaya):

You can do this two different ways. You can plug in 2 into each equation and see which gives you a true statement. 2=2 or 0=0 or 1=1 or something like is a true statement 2=5 or -6=1 is not a true statement or you could solve each of them

OpenStudy (anonymous):

–4x – 6 + x = –2x + x – 10 -4(2)-6+2=-2(2)+2-10 -8-6+2=-4+2-10 -12=-12